OpenStudy (anonymous):

\[f(x)=sin x = \sum_{n=0}^{\infty} \frac{f^n (\frac{\pi}{2})}{2n!} (x-\frac{\pi}{2})^{2n}\] It's an alternating series, how do I show that in my sum equation

OpenStudy (anonymous):

\[(-1)^n\]?

OpenStudy (anonymous):

a= pi/2

OpenStudy (anonymous):

\[ \sum _{n=0}^\infty \frac{(-1)^n \left(x-\frac{\pi }{2}\right)^{2 n}}{(2 n)!} \]

OpenStudy (anonymous):

yay!...so (-1)^n for alternating series

OpenStudy (anonymous):

wait lets compare \[\sum _{n=0}^\infty \frac{(-1)^n \left(x-\frac{\pi }{2}\right)^{2 n}}{(2 n)!}\] \[\sum_{n=0}^{\infty} \frac{f^n (\frac{\pi}{2})}{2n!} (x-\frac{\pi}{2})^{2n}\] so what happened to f^n (pi/2)...is that not necessary to write?

OpenStudy (anonymous):

Here are the first 21 derivatives of sin(x) \[ \begin{array}{cc} 0 & \sin (x) \\ 1 & \cos (x) \\ 2 & -\sin (x) \\ 3 & -\cos (x) \\ 4 & \sin (x) \\ 5 & \cos (x) \\ 6 & -\sin (x) \\ 7 & -\cos (x) \\ 8 & \sin (x) \\ 9 & \cos (x) \\ 10 & -\sin (x) \\ 11 & -\cos (x) \\ 12 & \sin (x) \\ 13 & \cos (x) \\ 14 & -\sin (x) \\ 15 & -\cos (x) \\ 16 & \sin (x) \\ 17 & \cos (x) \\ 18 & -\sin (x) \\ 19 & -\cos (x) \\ 20 & \sin (x) \\ \end{array} \]

OpenStudy (anonymous):

Those derivatives evaluated at \(\pi/2 \) \[ \begin{array}{cc} 0 & 1 \\ 1 & 0 \\ 2 & -1 \\ 3 & 0 \\ 4 & 1 \\ 5 & 0 \\ 6 & -1 \\ 7 & 0 \\ 8 & 1 \\ 9 & 0 \\ 10 & -1 \\ 11 & 0 \\ 12 & 1 \\ 13 & 0 \\ 14 & -1 \\ 15 & 0 \\ 16 & 1 \\ 17 & 0 \\ 18 & -1 \\ 19 & 0 \\ 20 & 1 \\ \end{array} \]

OpenStudy (anonymous):

yep I made a table n f^n x f^n (pi/2) 0 sinx 1 1 cosx 0 2 -sinx -1 3 -cosx 0 4 sinx 1

OpenStudy (anonymous):

but in my final summation notation do I not need to say f^(n) (pi/2)

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok so we did this problem together yesterday and the final answer then is \[f(x)=\frac{1}{x}=\sum_{n=0}^{\infty} \frac{(x+3)^n}{n!}\]

OpenStudy (anonymous):

for a=-3

OpenStudy (anonymous):

and what?

OpenStudy (anonymous):

oh I see you were trying to explain the series to me that's why you left the f^n (-3) in the summation notation. Makes sense now